A machinist's lathe was purchased 8 years ago and, with installation and setup, has an initial cost of $270,000. It will have a salvage value of 23,900 6 years from now. Assuming straight-line depreciation is a reasonable model, determine the current book value.(Use 3 significant figures for all calculations in this question)
Answer #1

Annual depreciation will be:

Annual\, depreciation=\frac{Machine\, cost-Salvage\, value}{Machine\, life}

Annual\, depreciation=\frac{270,000-23,900}{14}

Annual\, depreciation=17,578.57

The current book value of machine would be:

$270,000 - ($17,578.57 × 8) = $129,371.43